5(w+3)+9=3(w-2)+6 this is the problem for my son's homework and I can not remember the formula for solving this type of equation? Can you help?
Answer by bam878s(77)

First distribute the 5 and 3 on both sides.
5(w+3) = 5w + 15
3(w-2) = 3w - 6
so, we have
(5w+15) + 9 = (3w-6) + 6.
5w+24 = 3w.
subtract 3w from both sides.
2w+24=0.
subtract 24 from both sides.
2w=-24
divide both sides by 2.
w=-12.